OpenStudy (anonymous):

How I solve the following 3 problems? a is equal to theta in all of them. 1. (cot a +1)(csc a - .5) = 0 2. sin (square root 3 a)/2 = - (square root 3) /2 3. 2 cos 3a + 1 =0

OpenStudy (anonymous):

For number 2 I got a = -1

OpenStudy (anonymous):

In the 1.), for the product to equal 0, either cota + 1 = 0 or csca - 5 = 0. Solve for a in each case and that will give you the values for in that question. For 2.), you first find where is the sine function equal to -sqrt(3) / 2. We know that happens when the brackets are equal to 4pi/3 or 5pi/3. So solve sqrt(3a)/2 = 4pi/3 or sqrt(3a)/2 = 5pi/3, and solve for a in each case to get all the values. In 3.) Rearrange the equation so that cos(3a) = -1/2. Now we know that cosine is negative in the second and third quadrants. We also know that cos(pi/3) = 1/2. So the value must be -1/2 at 2pi/3 and 4pi/3. Once again, equate 3a = 2pi/3 or 3a = 4pi/3 and solve for a in each case to get all values. @wedontneednoedukation

OpenStudy (anonymous):

Do you understand? @wedontneednoedukation

OpenStudy (anonymous):

For question 1. I would do cot a = -1 right? So tan -1 = a correct? And I would also do csc a = .5 right? so cos .5 = a right? and both of those would be the answer?

OpenStudy (anonymous):

None of this helped at all, you are terrible.